OpenStudy (anonymous):

The quotient of a number and two is increased by seven. The result is eleven. What is the number? 9 1 8 2

OpenStudy (anonymous):

4 duh

OpenStudy (bibby):

Prove it with out guess-check

OpenStudy (anonymous):

?

OpenStudy (bibby):

\[\frac{x}{2}+7=11\]

OpenStudy (anonymous):

8 ma bad 8/2 = 4 4+7 = 11

OpenStudy (anonymous):

what is it 11 or 2

OpenStudy (bibby):

It's 8.

OpenStudy (anonymous):

and by checkin like bibby said, everything times 2, to isolate x 2*7 and 2*11 x + 14 = 22 -14 -14 x = 11

OpenStudy (anonymous):

9,1,8,2 what one

OpenStudy (anonymous):

8 .. -_-

OpenStudy (anonymous):

What is the solution to x - 9 > -15? x < -6 x > -6 x < -24 x > -24

OpenStudy (anonymous):

B

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

x>-6

OpenStudy (anonymous):

@Ello_M8!

OpenStudy (anonymous):

yus..

OpenStudy (anonymous):

How would you graph the solution to 3x ≤ -15? Draw an open circle on -5 and shade to the left. Draw an open circle on -5 and shade to the right. Draw a closed circle on -5 and shade to the left. Draw a closed circle on -5 and shade to the right.

OpenStudy (bibby):

Ask another question

OpenStudy (bibby):

Oh you already have

OpenStudy (anonymous):

C duhhh

OpenStudy (anonymous):

Solve d - (-2) = 19 for d. d =

OpenStudy (anonymous):

@Ello_M8! @bibby

OpenStudy (anonymous):

17